Circular No. 7690

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001du IN NGC 1365 R. Evans, Hazelbrook, N.S.W., reports his visual discovery with a 0.31-

m reflector on Aug. 24.7 and 25.7 UT of an apparent supernova (mag ∼ 14.0) located ∼ 90′′ west and 10′′ south of the nucleus of NGC 1365, and ∼ 20′′ west of a foreground star of mag ∼ 15.5. Evans adds that the new object was not seen on Aug. 23.7 under poorer conditions. A confirming CCD image obtained by G. Bock and P. Marples (Edens Landing, Queensland; 0.25-m Schmidt-Cassegrain reflector) shows SN 2001du at mag ∼ 14.1 on Aug. 24.79, and Bock provides the following approximate position for the new object: α = 3^{h}33^{m}28^{s}.7, δ = −36^{o}08′32′′ (equinox 2000.0). G. J. Garradd, Loomberah, N.S.W., adds that SN 2001du, which is not visible to mag ∼ 18.5 on an image taken on Aug. 15.76, appears at the western end of the galaxy’s bar and is superimposed on some H II regions. M. Salvo and B. Schmidt, Australian National University (ANU); M. Ashley and A. Phillips, University of New South Wales (UNSW); and C. Stubbs, University of Washington, report that SN 2001du is present at V = 14.8 ± 0.5 on a prediscovery CCD image taken with the UNSW Automated Patrol Telescope at Siding Spring on Aug. 23.81. It also appears at V = 15.2 ± 0.1 on images taken by P. Francis (ANU) with the Mt. Stromlo 1.9-m telescope on Aug. 25.65. Schmidt adds that the new object is not visible on a Digitized Sky Survey (DSS) image. A. Hale, Las Cruces, NM, reports that a visual observation of SN 2001du on Aug. 26.47 yields mag 14.2 (agreeing that nothing is present at that location on the red DSS image).

SUPERNOVA 2001dv IN MCG +09-27-71 T. Puckett, Mountain Town, GA; and D. George, Ottawa, ON, re-

port the discovery of an apparent supernova (mag 18.5) on an unfiltered CCD frame (limiting mag 20.0) taken with the Puckett Observatory 0.60m automated supernova patrol telescope on Aug. 23.15 UT. SN 2001dv is located at α = 16^{h}44^{m}56^{s}.04, δ = +51^{o}15′25′′.8 (equinox 2000.0), which is 3′′.0 east and 5′′.8 south of the center of MCG +09-27-71. The new object was confirmed on CCD frames taken on Aug. 24.19 with a 0.40-m reflector by J. Newton, Osoyoos, BC. The object was also confirmed on unfiltered images taken by Puckett and M. Marcus on Aug. 25.19. SN 2001dv is not present on images taken by Puckett on 1999 June 12, 1999 Aug. 20, 2000 Mar. 8, Apr. 16, and May 21 (limiting mag ∼ 20.5), neither is it present on Palomar Sky Survey images taken on 1992 June 3, 1990 Apr. 26 (limiting mag ∼ 21.0), and 1954 June 29 (limiting mag ∼ 19.7).

2001 August 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green